OpenStudy (anonymous):

Quadrilateral ABCD with vertices A(4, 3), B(4, -2), C(-4, -2) and D(-4, 3) is a rectangle, find the length of the diagonals.

jimthompson5910 (jim_thompson5910):

use the distance formula to find the distance from A to C

jimthompson5910 (jim_thompson5910):

d = sqrt( (x2 - x1)^2 + (y2 - y1)^2 ) d = sqrt( (-4 - 4)^2 + (-2 - 3)^2 ) d = sqrt( (-8)^2 + (-5)^2 ) d = sqrt( 64 + 25 ) d = sqrt( 89 ) So that's the length of AC, which is one of the diagonals

jimthompson5910 (jim_thompson5910):

I'll let you do the other diagonal

OpenStudy (anonymous):

That made no sense to me..

jimthompson5910 (jim_thompson5910):

have you heard of the distance formula before?

OpenStudy (anonymous):

I dunno what it is

jimthompson5910 (jim_thompson5910):

alright, it's just a formula to calculate the distance between any two points

jimthompson5910 (jim_thompson5910):

if you never seen it before, then I suggest looking at this page http://www.purplemath.com/modules/distform.htm it gives a good visual tutorial on how it works